# Decipher-MR: A Vision-Language Foundation Model for 3D MRI Representations

### v1

@@ 4.6 Pretraining Implementation Details
During the initial stage of vision encoder pretraining, the model was trained for 100 epochs with a batch size of 64 and a peak learning rate of 0.002. We employed a cosine learning rate scheduler with a linear warmup over the first 10 epochs, a teacher momentum consine schedule from 0.992 to 1, and a weight decay cosine schedule from 0.04 to 0.4. For the text encoder, we initialized it from PubMedBERT and fine-tuned it on our report data for 100 epochs with a learning rate of 5e-5 and an MLM probability of 0.15.
@@ 4.8 Classification and Regression
For each task, we applied a 3-layer MLP with 768-dimensional hidden layers, 20% dropout, and ReLU activations to the CLS token of the backbone ViT, serving as the classification or regression head. All evaluations and comparisons followed a simple probing setup, with the weights of the pretrained vision backbone frozen.
@@
To ensure reproducibility, we used 50 random train/validation/test splits for all experiments. We report the average AUROC for classification tasks and the average mean absolute error (MAE) for age regression.
@@ 4.9 Cross-Modal Retrieval
Using the validation split from our pretraining dataset, we performed paired report-to-image retrieval, where both full reports and conclusion sections were used as queries. Retrieval success was quantified as the percentage of cases in which the ground-truth MRI image appeared within the top 5, 10 and 20 results.
@@ 4.10 Segmentation
For the decoder, we adapted architectures from the SegResNet framework[46], a ResNet-based segmentation model. The decoder was modified to leverage patch embeddings from the final and three intermediate layers of the encoder, using skip connections between encoder and decoder stages to preserve spatial detail and support segmentation of complex anatomical structures.
All images and masks were resampled to standardized voxel resolutions prior to training. In addition to resampling to median voxel spacing as in nnUNet, we also experimented with multiple resolution levels across each dataset—including the 25th percentile, 10th percentile, and smallest voxel spacing (i.e., highest resolution)—to assess the impact of input resolution on performance.
@@ 7 Competing Interests
The authors declare no competing interests.

### v2

@@ 4.8 Pretraining Implementation Details
During the initial stage of vision encoder pretraining, the model was trained for 100 epochs with a batch size of 64 and a peak learning rate of 0.002. We employed a cosine learning rate scheduler with a linear warmup over the first 10 epochs, a teacher momentum consine schedule from 0.992 to 1, and a weight decay cosine schedule from 0.04 to 0.4, and, for iBOT, a mask ratio randomly sampled in 0.1-0.5. For the text encoder, we initialized it from PubMedBERT and fine-tuned it on our report data for 100 epochs with a learning rate of 5e-5 and an MLM mask ratio of 0.15.
@@ 4.9 Baselines
MRI-CORE[10]: A 2D MRI-specific foundation model pretrained on more than 6 million slices from over 110,000 MRI volumes across 18 body locations.
BrainSegFounder[7]: A 3D brain MRI segmentation foundation model pretrained in a multi-stage self-supervised framework on large-scale brain MRI data from 41,400 participants. We use public Swin encoder as a feature extractor for our evaluation tasks.
@@ 4.10 Classification and Regression
For each task, we applied a 3-layer MLP with 768-dimensional hidden layers, 20% dropout, and ReLU activations (1.2M parameters) to the CLS embedding of the backbone ViT, serving as the classification or regression head. All evaluations and comparisons followed a simple probing setup, with the weights of the pretrained vision backbone frozen. For ACDC, we extracted embeddings from both the ED (end-diastolic) and ES ((end-systolic) phases of the cine-MRI and concatenated the resulting embeddings as input to the MLP. A consistent learning rate of 1e-4 with plateau decay was applied across all tasks, without intense hyper-parameter tuning. Classification performance and model comparisons were generally robust to the choice of learning rate, given the simplicity of the probing setup on the extracted embeddings. (Supplementary Table 7)
@@
To ensure reproducibility, we used 50 random train/validation/test splits for all experiments. We report the average AUROC for classification tasks and the average mean absolute error (MAE) for age regression. Additionally, we performed bias analyses by training and testing separately on male and female cohorts for brain age, Alzheimer’s disease (793 females; 993 males), and body-region prediction (2,630 females; 1,790 males) to evaluate cross-group generalizability.
@@ 4.11 Cross-Modal Retrieval
Using the held-out test set from our pretraining dataset, we performed paired report-to-image retrieval, using both full radiology reports and conclusion sections as text queries. Retrieval performance was quantified as the proportion of cases in which the exact ground-truth MRI image (among approximately 25,000 MR images from 2,500 studies) appeared within the top 5, top 10, or top 20 retrieved results.
@@ 4.12 Segmentation
For the decoder, we adapted architectures from the SegResNet framework[28], a ResNet-based segmentation model. The decoder was modified to leverage patch embeddings from the final and three intermediate layers of the encoder, using skip connections between encoder and decoder stages to preserve spatial detail and support segmentation of complex anatomical structures. Overall, the resulting network contains approximately 7.6 million parameters.
All images and masks were resampled to standardized voxel resolutions prior to training. In addition to resampling to median voxel spacing as in nnUNet, we also experimented with multiple resolution levels across each dataset—including the 25th percentile, 10th percentile, and smallest voxel spacing (i.e., highest resolution)—to assess the impact of input resolution on performance. A learning rate of 1e-3 was used across all tasks. We generally found that a higher learning rate improved performance in all segmentation tasks, although excessively large values led to training instability and collapse.
@@ 7 Competing Interests
All authors are employees of GEHC. The authors declare no other competing interests.